# Categorical Syllogisms Revisited: A Review of the Logical Reasoning Abilities of LLMs for Analyzing Categorical Syllogisms

### v1

@@ 4.2.1 Setups
We conduct our experiments using OpenAI’s GPT models, as they are commonly used large language models with compelling performances on a variety of inference tasks OpenAI et al. (2024). All our experiments are done using GPT-4 and GPT-4o. We use the same set of datasets that we assess in Section 3.3. The details of these datasets are provided in Section 3.3.2.
Used Prompts.
@@ 4.2.2 Results
We visualize the error rate of GPT-4 and GPT-4o on a complete of 256 configurations of syllogisms from the Reasoning dataset in Figures 1(a) and 1(b). The error rate in the SylloFigure and Avicenna datasets are reported in Figure 1(c). We also report the total accuracy of validity judgment in Table 6 for reference purposes.
We observe the following trends. (1) Comparing Figure 1(a) with Figure 1(b), we observe different patterns for the configurations of syllogism that LLMs fail. For example, GPT-4 nearly has no errors when two premises are in AA format, while GPT-4o makes even more than half of the mistakes for AAI-3 and AAI-4. However, GPT-4o performs better than GPT-4 for configurations that GPT-4 has 0% accuracy. (2) For two datasets with human written text, GPT-4 seems to have more stable performances compared to GPT-4o, i.e., the error rate in Figure 1(c) is roughly the same for AAA-1, AAI-1, and AII-1. (3) We observe that for the same configuration, LLMs generally have a higher error rate in human-generated SylloFigure and Avicenna datasets (Figure 1(c)), compared to the template-based Reasoning dataset (Figures 1(a) and 1(b)). It seems to suggest that translating the syllogisms to the standard form is the bottleneck for LLMs to behave well, as the only difference that the Reasoning dataset has is the expressed way of the premises and conclusions. The underlying ability required to infer remains unchanged: if LLMs can translate the ordinary text into the standard format, then it should work well. This observation also aligns with the challenges of the logicians’ approach for analyzing syllogisms: as discussed in Section 2.1, the most difficult part is translating the propositions – once the mood and the figure are determined, then checking the validity of the syllogism is trivial.
@@ 4.3 Ambiguity of Natural Language
Our observation is that translating into standard propositions would be the most challenging part for LLMs and thus cause errors. In this section, we take a closer look at the types of translation errors LLMs make, especially around quantifiers.
We visualize in Figure 2 the percentage of correct propositions within each predicted proposition type in the SylloFigure dataset. We observe that in general, the LLMs have a much higher accuracy in recognizing the “some” quantifier, although it sometimes mixes the particular negative type (O) with the particular affirmative type (I). We also observe LLMs tend to confuse universal affirmative (A) with particular affirmative (I): among 1,546 propositions that are predicted as universal affirmative type (A), 33.2% should be particular affirmative (I). This phenomenon is related to the interpretation of singular propositions (discussed in Appendix A) and is also partially due to the fact that the singular propositions take a huge portion in the SylloFigure dataset (shown in Table 5).
We shall point out that analyzing the sensitivity of quantifiers by LLMs is not entirely new in computation linguistics. One representative work is Cui et al. (2022), where the authors rely on generalized quantifier theory to quantify their contribution to the errors of NLU models. There is a most recent work by Madusanka et al. (2023) that tries to investigate how different generalized quantifiers affect LLMs by employing a textual entailment problem. Ando et al. (2023) also suggest the importance of differentiating the problems of interpreting quantifiers and negations from performing logical inferences. In this work, we hold the same standpoints that the comprehension of quantifiers greatly affects the model performances and future models should enhance their abilities to analyze quantifiers. Compared to these prior studies, we present a more complete and comprehensive analysis of quantifiers in a specific syllogism setting.
@@ 5.1 Suggestions for Future Datasets
As shown in Table 4, many syllogism datasets lack certain kinds of annotations, thus causing troubles when we try to assess the coverage of language variations in Section 3.3. We notice that in their dataset descriptions, especially datasets that make use of templates, many annotations are actually generated during the dataset construction process (for example, blocks marked with  in Table 4). We suggest researchers could consider releasing these annotations from intermediate steps to promote a more accurate assessment of the properties of their datasets.
@@ 6 Conclusion
This work tries to address the question of whether the current proposed benchmarks could evaluate logical reasoning abilities accurately and thoroughly. We choose categorical syllogism as our main focus, since this logical system has been extensively studied by the logicians and has many nice properties, such as a finite number of possible cases, and automated ways of solving it. Categorical syllogism is also not arguably the most basic building block for any other more complex reasonings. We draw the inspirations of how logicians analyze categorical syllogisms and construct a list of variations that should be covered by the benchmarks. Our results show that there is no single dataset that properly covers all possible situations. We also summarize the current progress made in judging the validity of the categorical syllogisms. Our findings highlight the importance of correctly interpreting different quantifiers. Finally, we provide a discussion of several points that might be worth considering when researchers plan on the future release of categorical syllogism datasets.
@@
In this work, we mainly focus on analyzing the existing benchmarks of categorical syllogisms. Among 6 syllogism datasets listed in Table 4, we are only able to assess 3, as the left others are not publicly released. Also, we use GPT-4 as an annotation tool instead of human annotators for making up the missing annotations, such as mood, figure, and forms of phraseology. Although we have taken steps to control the quality of these annotations (as discussed in Section 3.3.1), it is inevitable that there are errors.
@@ Appendix A Handling Special Cases When Analyzing Categorical Syllogisms
A valid syllogism must have exactly three terms. When more than three terms seem to be involved in an argument of apparently syllogistic form, we may need to reduce the number of terms to three, by either eliminating synonyms, or eliminating class components Copi et al. (2019).
@@
In real life, we normally do not make explicit mention of all the premises required to support a given conclusion, especially when the premises are obvious or noncontroversial. A syllogism with an unstated premise is called an enthymeme Kelley (2013). A sorites is defined as a chain of categorical syllogisms in which the intermediate conclusions have been left out Hurley and Watson (2018). The standard treatment for analyzing sorites is to first make its intermediate conclusions or steps explicit, then test the validity of obtained syllogisms separately.

### v2

@@ 4.2.1 Setups
We conduct our experiments using OpenAI’s GPT models, as they are commonly used large language models with compelling performance on a variety of inference tasks OpenAI et al. (2024). All our experiments are done using GPT-4 and GPT-4o. We use the same set of datasets that we assess in Section 3.3. The details of these datasets are provided in Section 3.3.2.
Prompts Used.
@@ 4.2.2 Results
We visualize the error rate of GPT-4 and GPT-4o on the complete 256 configurations of syllogisms from the Reasoning dataset in Figures 1(a) and 1(b). The error rate in the SylloFigure and Avicenna datasets are reported in Figure 1(c). We also report the total accuracy of validity judgment in Table 6 for reference purposes.
We observe the following trends. (1) Comparing Figure 1(a) with Figure 1(b), we observe different patterns for the configurations of syllogism that LLMs fail. For example, GPT-4 nearly has no errors when two premises are in AA format, while GPT-4o makes even more than half of the mistakes for AAI-3 and AAI-4. However, GPT-4o performs better than GPT-4 for configurations that GPT-4 has 0% accuracy. (2) For two datasets with human-written text, GPT-4 seems to have more stable performance compared to GPT-4o, i.e., the error rate in Figure 1(c) is roughly the same for AAA-1, AAI-1, and AII-1. (3) We observe that for the same configuration, LLMs generally have a higher error rate in human-generated SylloFigure and Avicenna datasets (Figure 1(c)), compared to the template-based Reasoning dataset (Figures 1(a) and 1(b)). It seems to suggest that translating the syllogisms to the standard form is the bottleneck for LLMs to behave well, as the only difference that the Reasoning dataset has is the expressed way of the premises and conclusions. The underlying ability required to infer remains unchanged: if LLMs can translate ordinary text into the standard format, then it should work well. This observation also aligns with the challenges of the logicians’ approach for analyzing syllogisms: as discussed in Section 2.1, the most difficult part is translating the propositions – once the mood and the figure are determined, then checking the validity of the syllogism is trivial.
@@ 4.3 Ambiguity of Natural Language
Our observation is that translating into standard propositions is the most challenging part for LLMs and thus causes errors. In this section, we take a closer look at the types of translation errors LLMs make, especially around quantifiers.
We visualize in Figure 2 the percentage of correct propositions within each predicted proposition type in the SylloFigure dataset. We observe that in general, the LLMs have a much higher accuracy in recognizing the “some” quantifier, although it sometimes mixes the particular negative type (O) with the particular affirmative type (I). We also observe LLMs tend to confuse universal affirmative (A) with particular affirmative (I): among 1,546 propositions that are predicted as universal affirmative type (A), 33.2% should be particular affirmative (I). This phenomenon is related to the interpretation of singular propositions (discussed in Appendix A) and is also partially due to the fact that singular propositions represent a huge portion of the SylloFigure dataset (shown in Table 5).
We shall point out that analyzing the sensitivity of quantifiers by LLMs is not entirely new in computation linguistics. One representative work is Cui et al. (2022), where the authors rely on generalized quantifier theory to quantify their contribution to the errors of NLU models. There is a recent work by Madusanka et al. (2023) that tries to investigate how different generalized quantifiers affect LLMs by employing a textual entailment problem. Ando et al. (2023) also suggest the importance of differentiating the problems of interpreting quantifiers and negations from performing logical inferences. In this work, we hold the same standpoints that the comprehension of quantifiers greatly affects the model performance and future models should enhance their abilities to analyze quantifiers. Compared to these prior studies, we present a more complete and comprehensive analysis of quantifiers in a specific syllogism setting.
@@ 5.1 Suggestions for Future Datasets
As shown in Table 4, many syllogism datasets lack certain kinds of annotations, thus causing trouble when we try to assess the coverage of language variations in Section 3.3. We notice that in their dataset descriptions, especially datasets that make use of templates, many annotations are actually generated during the dataset construction process (for example, blocks marked with  in Table 4). We suggest researchers consider releasing these annotations from intermediate steps to promote a more accurate assessment of the properties of their datasets.
@@ 6 Conclusion
This work tries to address the question of whether current proposed benchmarks can evaluate logical reasoning abilities accurately and thoroughly. We choose categorical syllogism as our main focus, since this logical system has been extensively studied by logicians and has many nice properties, such as a finite number of possible cases, and automated ways of solving it. A categorical syllogism is also arguably the most basic building block for any other more complex reasonings. We draw the inspirations from how logicians analyze categorical syllogisms and construct a list of variations that should be covered by benchmarks. Our results show that there is no single dataset that properly covers all possible situations. We also summarize the current progress made in judging the validity of the categorical syllogisms. Our findings highlight the importance of correctly interpreting different quantifiers. Finally, we provide a discussion of several points that might be worth considering when researchers plan on the future release of categorical syllogism datasets.
@@
In this work, we mainly focus on analyzing the existing benchmarks of categorical syllogisms. Among 6 syllogism datasets listed in Table 4, we are only able to assess 3, as others are not publicly released. Also, we use GPT-4 as an annotation tool instead of human annotators to generate the missing annotations, such as mood, figure, and forms of phraseology. Although we have taken steps to control the quality of these annotations (as discussed in Section 3.3.1), it is inevitable that there are errors.
@@ Appendix A Handling Special Cases When Analyzing Categorical Syllogisms
A valid syllogism must have exactly three terms. When more than three terms seem to be involved in an argument of apparently syllogistic form, we may need to reduce the number of terms to three, by either eliminating synonyms or eliminating class components Copi et al. (2019).
@@
In real life, we normally do not make explicit mention of all the premises required to support a given conclusion, especially when the premises are obvious or noncontroversial. A syllogism with an unstated premise is called an enthymeme Kelley (2013). Sorites are defined as a chain of categorical syllogisms in which the intermediate conclusions have been left out Hurley and Watson (2018). The standard treatment for analyzing sorites is to first make their intermediate conclusions or steps explicit, then test the validity of obtained syllogisms separately.